# Decoupled-Value Attention for Prior-Data Fitted Networks: GP Inference for Physical Equations

### v1

@@ Abstract
Prior-data fitted networks (PFNs) are a promising alternative to time-consuming Gaussian Process (GP) inference for creating fast surrogates of physical systems. PFN reduces the computational burden of GP-training by replacing Bayesian inference in GP with a single forward pass of a learned prediction model. However, with standard Transformer attention, PFNs show limited effectiveness on high-dimensional regression tasks. We introduce Decoupled-Value Attention (DVA)– motivated by the GP property that the function space is fully characterized by the kernel over inputs and the predictive mean is a weighted sum of training targets. DVA computes similarities from inputs only and propagates labels solely through values. Thus, the proposed DVA mirrors the Gaussian-process update while remaining kernel-free. We demonstrate that the crucial factor for scaling PFNs is the attention rule rather than the architecture itself. Specifically, our results demonstrate that (a) localized attention consistently reduces out-of-sample validation loss in PFNs across different dimensional settings, with validation loss reduced by more than 50% in five- and ten-dimensional cases, and (b) the role of attention is more decisive than the choice of backbone architecture, showing that CNN-based PFNs can perform at par with their Transformer-based counterparts. The proposed PFNs provide 64-dimensional power flow equation approximations with a mean absolute error of the order of $10^{-3}$, while being over $80\times$ faster than exact GP inference. Code: https://github.com/PSquare-Lab/DVA-PFN
@@ 1 Introduction
Bayesian inference provides a powerful framework for reasoning under uncertainty, with methods like Gaussian Processes (GPs) offering well-calibrated predictions and principled uncertainty estimates (Williams & Rasmussen, 2006). However, the practical application of these methods is often hindered by the heavy computational burden of learning kernel hyperparameters. For example, exact GP inference scales cubically with the number of data points, making its deployment infeasible for large datasets or problems requiring repeated training. Consider a physical system where a surrogate GP is chosen due to its uncertainty estimates and differentiable closed-form expressions. However, the underlying input dataset and configuration changes frequently, and the surrogate is supposed to work for these new, previously unseen variations. For example, changes in underlying physical networks for power grids Tan et al. (2025). In such conditions, GP needs to be trained repeatedly, incurring significant computing cost, each time the dataset changes.
@@
To address this, Prior-Data Fitted Networks (PFNs) have emerged as a method (Müller et al., 2022) that uses large-scale pre-training to approximate the Bayesian posterior predictive in a single forward pass. Note that unlike sparse GP approximations Daskalakis et al. (2022), PFNs eliminate kernel parameter training step. Although Low-rank approximations reduce GP cost to $\mathcal{O}(nm^{2})$, where $m$ is a user-defined parameter, PFNs need only a forward pass at deployment. This advantage grows when multiple GPs must be learned, as training $K$ GPs scales to $\mathcal{O}(Knm^{2})$, with each $m$ requiring tuning. PFNs avoid these issues by directly predicting the posterior distribution in one step in a forward pass of the trained network. However, PFNs face scaling and bias issues in problems with high input dimensions due to their joint input–output embedding strategy Müller et al. (2022); Hollmann et al. (2025); Wang et al. (2025); Nagler (2023). Attention over concatenated $(\mathbf{x},\mathbf{y})$ embeddings, as done in PFNs, degrade locality and similarity measures as input dimension grows Further, they are almost exclusively built using Transformer architectures, which have high memory requirements. These challenges in existing PFFs motivate this work.
@@
In this work, we propose Decoupled-Value Attention (DVA), an input localized attention mechanism to scale PFNs with different architectures. We provide evidence that the attention mechanism is the primary driver of PFN performance, and it can be built using different architectures such as Convolution Neural Networks (CNNs), and Transformers. The proposed DVA computes attention affinities (queries and keys) purely from the input space, while propagating information from the output space exclusively through the values. This aligns directly with the functional-space view of a GP, where the influence of training outputs $y_{i}$ on a test prediction is weighted by the similarity of their corresponding inputs $x_{i}$ Williams & Rasmussen (2006). This is a significant deviation from the standard attention mechanism applied in existing PFN works where affinities are calculated from a concatenated input-output vector Müller et al. (2022); Hollmann et al. (2022). This, combining inputs and outputs, increases the computational load, reducing PFNs ability to learn when the dimensions of input space grow. We note the observation made by Nagler (2023) that the convergence of PFNs is due to the attention mechanism, while bias [^1] Here, we use the definition of bias as the difference between the parametrized PPD and the true PPD, which converges to zero as the number of samples increases Nagler (2023). Consequently, a decrease in NLL reflects a reduction in this bias. is a function of architecture choice. More importantly, it argues that a post-hoc localization mechanism is needed to reduce bias. We show that using the proposed localized attention, the reduction in PFN validation loss is consistent across architectures and exceeds the bias variation caused by the architectures themselves. Experimental studies show that DVA performs better than standard Vanilla Attention (VA) used in PFN literature, across dimensions and architectures. Our main contributions are:
@@
  • A Localized Attention Mechanism for GP-PFNs: We introduce DVA, which explicitly enforces input-only localization and reduces difference between predicted and true posterior distributions in PFN training by more than 50% for the inputs of 5D and 10D [^2] We use D to indicate dimension; for example, ND means N-dimensional.. This design leads to substantially lower validation loss and improves predictive performance on high-dimensional regression tasks compared to standard PFN attention, without requiring additional data or compute resources.
@@
  • Attention is More Important than Architecture: We show that PFNs can also be constructed using CNN as backbone architecture, and with DVA, the choice of backbone architecture becomes secondary. This confirms that the attention mechanism is the primary driver of bias reduction. The proposed CNN-DVA based PFN achieves accuracy comparable to a Transformer-DVA based PFN across input dimensions upto 64D. Overall, changes in attention produce a more pronounced reduction in validation loss and predicted error than changes in the backbone architecture.
@@
  • Scaling PFNs to High-Dimensional Learning Problems: Standard PFNs with joint input–output attention fail to generalize beyond $\sim$10 input dimensions (10D), saturating at high validation loss. In contrast, DVA enables successful inference up to 64D and on power flow learning task in 64D, CNN+DVA achieves Mean Squared Error of order $10^{-5}$ even with 50% load uncertainty levels, and Mean Absolute Error on the order of $10^{-3}$ – at $\mathbf{80\times}$ the speed of exact GP inference.
@@
Positioning: We want to highlight that our goal is not to claim a novel, general-purpose attention mechanism. Rather, DVA is a specialized design intended to create scalable and robust PFNs via localization and emulation of GP inference. We also note that there are many efficient attention mechanisms, including linearized kernels Katharopoulos et al. (2020), Nyström approximations Xiong et al. (2021), random feature expansions Choromanski et al. (2021), and cross-kernel attention Wang & Others (2025), which are kernel-based. These attentions are designed to incorporate GP and kernel advantages into Transformer-based language and vision models, along with scaling approximations like Peng et al. (2021); Bui et al. (2025). In contrast, the proposed DVA is designed to develop scalable PFNs Hollmann et al. (2022) that can mimic GP inference for physical equations in particular. Further, our sole focus is not on scaling PFNs with Transformer-like Wang et al. (2025), instead a) highlight that attention is the critical component in PFNs over architectures and b) bias reduction in PFNs can be achieved via attention without post-hoc localization Nagler (2023). More importantly, DVA is intentionally designed to remain kernel-free because forcing a single kernel type can lead to significant model mismatch for physics problems. For instance, the functions governing AC power flow are best modeled by specialized kernels distinct from standard choices (Liu & Srikantha, 2022), and the optimal kernel can even change with operating conditions (Pareek & Nguyen, 2021). By learning a data-driven similarity metric, DVA remains flexible and robust, avoiding the need for manual kernel selection and tuning.
@@ 1.1 Related Works
Prior-data Fitted Networks: There are several works on PFNs (Hollmann et al., 2025; Wang et al., 2025; Nagler, 2023; Adriaensen et al., 2023; Li et al., 2023), most of which rely on the Transformer architecture (Vaswani et al., 2017), applying self-attention over concatenated $(x_{i},y_{i})$ embeddings. While this design has shown strong performance on certain tasks, it presents two key limitations that remain largely unaddressed. First, these works implicitly assume that the Transformer backbone is crucial to PFN success. Second—and more importantly—the standard attention mechanism does not scale well to high-dimensional problems: training becomes unstable, and performance deteriorates quickly as dimensionality increases Wang et al. (2025). Although Wang et al. (2025) introduced a Boosting-based method that splits the dataset into smaller subsets and trains an ensemble of PFNs, this was primarily intended to handle longer input sequences, not to address high-dimensional scaling issues or architectural dependence of PFNs.
Physical Equation Surrogates for Power Flow: Efficiently solving power flow equations is crucial for integrating renewable energy and electric vehicles Barry et al. (2022), a key area where machine learning can help mitigate climate change Rolnick et al. (2022). Faster analytical approximations of nonlinear alternating current power flow (ACPF) equations exist, but come at the cost of accuracy Molzahn et al. (2019). To address this, various ML models—including physics-informed methods—have been developed for ACPF learning and uncertainty quantification Chen et al. (2025). Among these, GPs have gained prominence for building explainable surrogates with closed-form predictions Tan et al. (2025). However, such modeling is extremely sensitive to GP kernels, as shown by Liu & Srikantha (2022) by showing that specialized kernels outperform standard options like squared-exponential or polynomial kernels Pareek & Nguyen (2021).
@@ 2.1 Gaussian Processes (GP)
GP is a non-parametric, probabilistic framework for modeling functions from a functional space perspective. Given data $(x_{i},y_{i})$, we assume $y_{i}=f(x_{i})+\varepsilon_{i}$, where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2}_{\varepsilon})$. For $N$ inputs $x=(x_{1},\dots,x_{N})$, the function values $\bm{f}(x)=[f(x_{1}),\dots,f(x_{N})]^{\top}$ follow a joint Gaussian distribution as $\bm{f}(x)\sim\mathcal{N}(\bm{m}(x),K(x,x^{\prime})),$ with mean function $m(x)$ and covariance matrix $K(x,x^{\prime})$. By definition, a GP is a collection of random variables such that any finite subset is jointly Gaussian, denoted $f\sim\mathcal{GP}(m(\cdot),k(\cdot,\cdot))$. The observation distribution is then $\mathbb{P}(y)\sim\mathcal{N}(m(x),K(x,x^{\prime})+\sigma_{\varepsilon}^{2}\mathbf{\rm I})$, with $\mathbf{\rm I}$ as identity matrix of appropriate size. Thus, given training data $y$ at $x$, the predictive distribution of $f^{\star}$ at a new input $x^{\star}$ is Gaussian with closed-form mean and covariance. The choice of kernel (covariance) function $k(x,x^{\prime})$ encodes prior assumptions about $f$, while hyperparameters are typically learned by maximizing the marginal log-likelihood. However, the closed-form of exact inference only works when the likelihood is Gaussian and inversion of kernel matrix presents training bottleneck. A key property of GPs, central to the design of DVA, is that the kernel $k(\cdot,\cdot)$ measures similarity solely between input data Williams & Rasmussen (2006).
@@ 2.2 Prior-Data Fitted Network (PFN)
PFNs (Müller et al., 2022) are neural predictors trained to approximate the posterior predictive distribution (PPD) of a Bayesian model in a single forward pass. Rather than fitting a single static dataset, a PFN is trained on multiple synthetic datasets– drawn from a prior over data-generating mechanisms. Given a prior distribution $p(\mathcal{D})$ over supervised learning tasks, PFNs repeatedly sample datasets $\mathcal{D}^{k}\cup\{(\mathbf{x}^{k},\mathbf{y}^{k})\}\sim p(\mathcal{D})$ for $k=1\dots K$ and train the model to minimize
@@
Here, $q_{\theta}(\cdot)$ represents the Transformer prediction. This procedure treats entire datasets $\mathcal{D}$’s as inputs and optimizes the model parameters $\theta$ to predict a held-out label conditioned on the remaining data. Thus, fitting the PPD without explicitly computing posteriors. Further, PFNs represent the output distribution using a discrete set of buckets (bins) for the target $\mathbf{y}$, essentially posing regression as a classification problem. After training, the PFN performs amortized Bayesian inference: given a new dataset $\mathcal{D}_{\text{train}}$ and query point $x_{\text{test}}$, it outputs $q_{\theta^{\star}}(y_{\text{test}}\mid x_{\text{test}},\mathcal{D}_{\text{train}})\approx p(y_{\text{test}}\mid x_{\text{test}},\mathcal{D}_{\text{train}})$ in a single forward pass, where $\theta^{\star}$ is optimal Transformer parameters Müller et al. (2022).

### v2

@@ Abstract
Prior-data fitted networks (PFNs) are a promising alternative to time-consuming Gaussian process (GP) inference for creating fast surrogates of physical systems. PFN reduces the computational burden of GP-training by replacing Bayesian inference in GP with a single forward pass of a learned prediction model. However, with standard Transformer attention, PFNs show limited effectiveness on high-dimensional regression tasks. We introduce Decoupled-Value Attention (DVA)– motivated by the GP property that the function space is fully characterized by the kernel over inputs and the predictive mean is a weighted sum of training targets. DVA computes similarities from inputs only and propagates labels solely through values. Thus, the proposed DVA mirrors the GP update while remaining kernel-free. We demonstrate that PFNs are backbone architecture invariant and the crucial factor for scaling PFNs is the attention rule rather than the architecture itself. Specifically, our results demonstrate that (a) localized attention consistently reduces out-of-sample validation loss in PFNs across different dimensional settings, with validation loss reduced by more than 50% in five- and ten-dimensional cases, and (b) the role of attention is more decisive than the choice of backbone architecture, showing that CNN, RNN and LSTM-based PFNs can perform at par with their Transformer-based counterparts. The proposed PFNs provide 64-dimensional power flow equation approximations with a mean absolute error of the order of $10^{-3}$, while being over $80\times$ faster than exact GP inference.
@@ 1 Introduction
Bayesian inference provides a powerful framework for reasoning under uncertainty, with methods like Gaussian processes (GPs) offering well-calibrated predictions and principled uncertainty estimates (williams2006gaussian). However, the practical application of these methods is often hindered by the heavy computational burden of learning kernel hyperparameters. For example, exact GP inference scales cubically with the number of data points, making its deployment infeasible for large datasets or problems requiring repeated training. Consider a physical system where a surrogate GP is chosen due to its uncertainty estimates and differentiable closed-form expressions. However, the underlying input dataset and configuration changes frequently, and the surrogate is supposed to work for these new, previously unseen variations. For example, changes in underlying physical networks for power grids tan2025gaussian. In such conditions, GP needs to be trained repeatedly, incurring significant computing cost, each time the dataset changes.
@@
To address this, Prior-Data Fitted Networks (PFNs) have emerged as a method (müller2024transformersbayesianinference) that uses large-scale pre-training to approximate the Bayesian posterior predictive in a single forward pass. Note that unlike sparse GP approximations lowrankGP, PFNs eliminate kernel parameter training step. Although Low-rank approximations reduce GP cost to $\mathcal{O}(nm^{2})$, where $m$ is a user-defined parameter, PFNs need only a forward pass at deployment. This advantage grows when multiple GPs must be learned, as training $K$ GPs scales to $\mathcal{O}(Knm^{2})$, with each $m$ requiring tuning. PFNs avoid these issues by directly predicting the posterior distribution in one step in a forward pass of the trained network. However, PFNs face scaling and bias issues in problems with high input dimensions due to their joint input–output embedding strategy müller2024transformersbayesianinference; hollmann2025tabpfn; wang25d; nagler23a. Attention over concatenated $(\mathbf{x},\mathbf{y})$ embeddings, as done in PFNs, degrade locality and similarity measures as input dimension grows Further, they are almost exclusively built using Transformer architectures, which have high memory requirements. These challenges in existing PFFs motivate this work.
@@
In this work, we propose Decoupled-Value Attention (DVA), an input localized attention mechanism to scale PFNs with different architectures. We provide evidence that the attention mechanism is the primary driver of PFN performance, and it can be built using different architectures Convolution Neural Networks (CNNs) as well along with Transformers. The proposed DVA computes attention affinities (queries and keys) purely from the input space, while propagating information from the output space exclusively through the values. This aligns directly with the functional-space view of a GP, where the influence of training outputs $y_{i}$ on a test prediction is weighted by the similarity of their corresponding inputs $x_{i}$ williams2006gaussian. This is a significant deviation from the standard attention mechanism applied in existing PFN works where affinities are calculated from a concatenated input-output vector müller2024transformersbayesianinference; hollmann2022pfn. This, combining inputs and outputs, increases the computational load, reducing PFNs ability to learn when the dimensions of input space grow. We note the observation made by nagler23a that the convergence of PFNs is due to the attention mechanism, while bias [^1] Here, we use the definition of bias as the difference between the parametrized PPD and the true PPD, with variance vanishing as the number of inference-time samples increases nagler23a. Since our experiments operate in this low-variance regime—as also evidenced by the negligible variance observed in our robustness results in Appendix D—any decrease in NLL necessarily reflects a reduction in this bias. is a function of architecture choice. More importantly, it argues that a post-hoc localization mechanism is needed to reduce bias. We provide both theoretical and empirical evidence that the proposed localized attention weights are proportional to the query–context distance during PFN inference. Further, the reduction in PFN validation loss is consistent across architectures. Experimental studies show that DVA performs better than standard Vanilla Attention (VA) used in PFN literature, across dimensions and architectures. Our main contributions are:
@@
  • A Localized Attention Mechanism for GP-PFNs: We introduce DVA, which explicitly enforces input-only localization and reduces difference between predicted and true posterior distributions in PFN training by more than 50% for the inputs of 5D and 10D [^2] We use D to indicate dimension; for example, ND means N-dimensional.. This design leads to substantially lower validation loss and improves predictive performance on high-dimensional regression tasks compared to standard PFN attention, without requiring additional data or compute resources. We prove the localization property of DVA both theoretically and empirically.
@@
  • Attention is More Important than Architecture: We show that PFNs can also be constructed using CNN as backbone architecture, and with DVA, the choice of backbone architecture becomes secondary. This confirms that the attention mechanism is the primary driver of bias reduction. The proposed CNN-DVA, RNN-DVA and LSTM-DVA based PFN achieves accuracy comparable to a Transformer-DVA based PFN across input dimensions upto 64D. Overall, changes in attention produce a more pronounced reduction in validation loss and predicted error than changes in the backbone architecture. We also show that DVA version of linear attention choromanski2022rethinkingattentionperformers works considerably better than VA, while being less effective than proposed softmax-DVA.
@@
  • Scaling PFNs to High-Dimensional Learning Problems: Standard PFNs with joint input–output attention fail to generalize beyond $\sim$10 input dimensions (10D), saturating at high validation loss. In contrast, DVA enables successful inference up to 64D on power flow learning task. The CNN+DVA achieves Mean Squared Error of order $10^{-5}$ even with 50% load uncertainty levels, and Mean Absolute Error on the order of $10^{-3}$ – at $\mathbf{80\times}$ the speed of exact GP inference.
@@
Positioning: We want to highlight that our goal is not to claim a novel, general-purpose attention mechanism. Rather, DVA is a specialized design intended to create scalable and robust PFNs via localization and emulation of GP inference. We also note that there are many efficient attention mechanisms, including linearized kernels katharopoulos2020transformers, Nyström approximations xiong2021nystromformer, random feature expansions choromanski2021rethinking, and cross-kernel attention wang2025crosskernel, which are kernel-based. These attentions are designed to incorporate GP and kernel advantages into Transformer-based language and vision models, along with scaling approximations like peng2021random; bui2025revisiting. In contrast, the proposed DVA is designed to develop scalable PFNs hollmann2022pfn that is suitable for physical equations in particular. Further, our sole focus is not on scaling PFNs with Transformer-like wang25d, instead a) highlight that attention is the critical component in PFNs over architectures and b) bias reduction in PFNs can be achieved via attention without post-hoc localization nagler23a. Since localization of attention for PFNs is the central idea of proposed work, we show that linear attention choromanski2022rethinkingattentionperformers for PFN in decoupled fashion improves performance—relative to coupled linear attention, although softmax-DVA achieves the lowest validation loss. More importantly, DVA is intentionally designed to remain kernel-free because forcing a single kernel type can lead to significant model mismatch for physics problems. For instance, the functions governing AC power flow are best modeled by specialized kernels distinct from standard choices (9734745), and the optimal kernel can even change with operating conditions (pareek2021framework). By learning a data-driven similarity metric, DVA remains flexible and robust, avoiding the need for manual kernel selection.
@@ 1.1 Related Works
Prior-data Fitted Networks: There are several works on PFNs (hollmann2025tabpfn; wang25d; nagler23a; adriaensen23lc; li23l), most of which rely on the Transformer architecture (vaswani2017attention), applying self-attention over concatenated $(x_{i},y_{i})$ embeddings. While this design has shown strong performance on certain tasks, it presents two key limitations that remain largely unaddressed. First, these works implicitly assume that the Transformer backbone is crucial to PFN success. Second—and more importantly—the standard attention mechanism does not scale well to high-dimensional problems: training becomes unstable, and performance deteriorates quickly as dimensionality increases wang25d. Although wang25d introduced a Boosting-based method that splits the dataset into smaller subsets and trains an ensemble of PFNs, this was primarily intended to handle longer input sequences, not to address high-dimensional scaling issues or architectural dependence of PFNs.
Physical Equation Surrogates for Power Flow: Efficiently solving power flow equations is crucial for integrating renewable energy and electric vehicles barry2022risk, a key area where machine learning can help mitigate climate change rolnick2022tackling. Faster analytical approximations of nonlinear alternating current power flow (ACPF) equations exist, but come at the cost of accuracy molzahnsurvey. To address this, various ML models—including physics-informed methods—have been developed for ACPF learning and uncertainty quantification ML_OPF_wiki. Among these, GPs have gained prominence for building explainable surrogates with closed-form predictions tan2025gaussian. However, such modeling is extremely sensitive to GP kernels, as shown by 9734745 by showing that specialized kernels outperform standard options like squared-exponential or polynomial kernels pareek2021framework.
@@ 2.1 Gaussian Processes (GP)
GP is a non-parametric, probabilistic framework for modeling functions from a functional space perspective. Given data $(x_{i},y_{i})$, we assume $y_{i}=f(x_{i})+\varepsilon_{i}$, where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2}_{\varepsilon})$. For $N$ inputs $x=(x_{1},\dots,x_{N})$, the function values $\bm{f}(x)=[f(x_{1}),\dots,f(x_{N})]^{\top}$ follow a joint Gaussian distribution as $\bm{f}(x)\sim\mathcal{N}(\bm{m}(x),K(x,x^{\prime})),$ with mean function $m(x)$ and covariance matrix $K(x,x^{\prime})$. By definition, a GP is a collection of random variables such that any finite subset is jointly Gaussian, denoted $f\sim\mathcal{GP}(m(\cdot),k(\cdot,\cdot))$. The observation distribution is then $\mathbb{P}(y)\sim\mathcal{N}(m(x),K(x,x^{\prime})+\sigma_{\varepsilon}^{2}\mathbf{\rm I})$, with $\mathbf{\rm I}$ as identity matrix of appropriate size. Thus, given training data $y$ at $x$, the predictive distribution of $f^{\star}$ at a new input $x^{\star}$ is Gaussian with closed-form mean and covariance. The choice of kernel (covariance) function $k(x,x^{\prime})$ encodes prior assumptions about $f$, while hyperparameters are typically learned by maximizing the marginal log-likelihood. However, the closed-form of exact inference only works when the likelihood is Gaussian and inversion of kernel matrix presents training bottleneck. A key property of GPs, central to the design of DVA, is that the kernel $k(\cdot,\cdot)$ measures similarity solely between input data williams2006gaussian.
@@ 2.2 Prior-Data Fitted Network (PFN)
PFNs (müller2024transformersbayesianinference) are neural predictors trained to approximate the posterior predictive distribution (PPD) of a Bayesian model in a single forward pass. Rather than fitting a single static dataset, a PFN is trained on multiple synthetic datasets– drawn from a prior over data-generating mechanisms. Given a prior distribution $p(\mathcal{D})$ over supervised learning tasks, PFNs repeatedly sample datasets $\mathcal{D}^{k}\cup\{(\mathbf{x}^{k},\mathbf{y}^{k})\}\sim p(\mathcal{D})$ for $k=1\dots K$ and train the model to minimize
@@
Here, $q_{\theta}(\cdot)$ represents the Transformer prediction. This procedure treats entire datasets $\mathcal{D}$’s as inputs and optimizes the model parameters $\theta$ to predict a held-out label conditioned on the remaining data. Thus, fitting the PPD without explicitly computing posteriors. Further, PFNs represent the output distribution using a discrete set of buckets (bins) for the target $\mathbf{y}$, essentially posing regression as a classification problem. After training, the PFN performs amortized Bayesian inference: given a new dataset $\mathcal{D}_{\text{train}}$ and query point $x_{\text{test}}$, it outputs $q_{\theta^{\star}}(y_{\text{test}}\mid x_{\text{test}},\mathcal{D}_{\text{train}})\approx p(y_{\text{test}}\mid x_{\text{test}},\mathcal{D}_{\text{train}})$ in a single forward pass, where $\theta^{\star}$ is optimal Transformer parameters müller2024transformersbayesianinference.